# Unleashing the Potential of Multi-modal Foundation Models and Video Diffusion for 4D Dynamic Physical Scene Simulation

### v1

@@ B.1 Optimization Details
In our experiments, we implement the differentiable MPM using Warp [12]. Besides, we use RAFT [36] to compute optical flow, which serves as a critical component for guiding the optimization process. For the real-world dataset, we segment the foreground points where input forces are applied using SAM2Point [11], ensuring precise identification of the regions impacted by force. Video generation is facilitated by CogVideoX [42], which leverages a text prompt (refer to Tab. 5) combined with a selected frame from the dataset to guide the synthesis process. All optimization tasks are performed on a single NVIDIA A800 GPU.
@@ B.2 Baselines
PhysGaussian/PhysDreamer/Physics3D: PhysGaussian utilizes a fixed set of manually defined material properties, including material type, density, and associated parameters, without further optimization. This setup is comparable to the conditions outlined in Sec. 4.3 (1). Changes to the material type or density, or further parameter optimization, can produce different simulation outcomes. To highlight the benefits of incorporating a large pre-trained visual foundation model for physics reasoning and parameter refinement, we apply the same input force to each scene and compare our results with PhysGaussian. While PhysDreamer and Physics3D include methods for parameter optimization, they are limited to the elastic material type. For consistency and fair comparison, we assign elastic as the material type in these methods while using the same parameters as PhysGaussian for other aspects.
@@ C.1 Human Evaluation
C.2 Additional Qualitative Results of Synthetic and Real-world Datasets
We provide additional visual comparisons between our method and baselines on both synthetic (see Fig. 7) and real-world (see Fig. 9) datasets. These results emphasize the effectiveness of our method in handling diverse material behaviors and complex scenarios. To further enhance understanding and visualization, we offer more results in videos, which are accessible through our project page: https://zhuomanliu.github.io/PhysFlow/.
Object $\mathcal{L}_{render}$ $\mathcal{L}_{sds}$ $\mathcal{L}_{flow}$ (Ours)
Droplet $\Delta_{\mu}=1.88\times 10^{4}$, $\Delta_{\kappa}=8.73\times 10^{4}$ $\Delta_{\mu}=1.22\times 10^{4}$, $\Delta_{\kappa}=1.26\times 10^{6}$ $\Delta_{\mu}=\mathbf{0.89}$, $\Delta_{\kappa}=\mathbf{0.3\times 10^{4}}$
Letter $\Delta_{\mu}=406.91$, $\Delta_{\kappa}=4.45\times 10^{4}$ $\Delta_{\mu}=605.24$, $\Delta_{\kappa}=2.27\times 10^{4}$ $\Delta_{\mu}=\mathbf{2.27}$, $\Delta_{\kappa}=\mathbf{0.7\times 10^{4}}$
Cream $\Delta_{\mu}=1.24\times 10^{6}$, $\Delta_{\kappa}=2.79\times 10^{7}$, $\Delta_{\mu}=1.26\times 10^{6}$, $\Delta_{\kappa}=2.83\times 10^{7}$, $\Delta_{\mu}=\mathbf{0.121\times 10^{5}}$, $\Delta_{\kappa}=\mathbf{5.3\times 10^{5}}$,
$\Delta_{\tau_{Y}}=2.87\times 10^{3}$, $\Delta_{\eta}=7.8$ $\Delta_{\tau_{Y}}=2.86\times 10^{3}$, $\Delta_{\eta}=7.8$ $\Delta_{\tau_{Y}}=\mathbf{1.73\times 10^{3}}$, $\Delta_{\eta}=\mathbf{7.5}$
Toothpaste $\Delta_{\mu}=4.21\times 10^{3}$, $\Delta_{\kappa}=0.82\times 10^{5}$, $\Delta_{\mu}=4.20\times 10^{3}$, $\Delta_{\kappa}=0.82\times 10^{5}$, $\Delta_{\mu}=\mathbf{0.07\times 10^{3}}$, $\Delta_{\kappa}=\mathbf{0.13\times 10^{5}}$,
$\Delta_{\tau_{Y}}=180$, $\Delta_{\eta}=584.14$ $\Delta_{\tau_{Y}}=180$, $\Delta_{\eta}=557.73$ $\Delta_{\tau_{Y}}=\mathbf{46}$, $\Delta_{\eta}=\mathbf{28.99}$
Torus $\Delta_{E}=7.02\times 10^{5}$, $\Delta_{\nu}=\mathbf{0.196}$ $\Delta_{E}=5.39\times 10^{5}$, $\Delta_{\nu}=0.212$ $\Delta_{E}=\mathbf{3.9\times 10^{4}}$, $\Delta_{\nu}=0.297$
Bird $\Delta_{E}=2.87\times 10^{5}$, $\Delta_{\nu}=0.221$ $\Delta_{E}=2.80\times 10^{5}$, $\Delta_{\nu}=0.197$ $\Delta_{E}=\mathbf{0.12\times 10^{5}}$, $\Delta_{\nu}=\mathbf{0.171}$
Playdoh $\Delta_{E}=1.55\times 10^{6}$, $\Delta_{\nu}=\mathbf{0.098}$, $\Delta_{E}=1.60\times 10^{6}$, $\Delta_{\nu}=\mathbf{0.098}$, $\Delta_{E}=\mathbf{0.21\times 10^{6}}$, $\Delta_{\nu}=\mathbf{0.098}$,
$\Delta_{\tau_{Y}}=1.48\times 10^{4}$ $\Delta_{\tau_{Y}}=1.48\times 10^{4}$ $\Delta_{\tau_{Y}}=\mathbf{41}$
Cat $\Delta_{E}=2.56\times 10^{6}$, $\Delta_{\nu}=\mathbf{0.102}$, $\Delta_{E}=2.61\times 10^{6}$, $\Delta_{\nu}=\mathbf{0.102}$, $\Delta_{E}=\mathbf{3.87\times 10^{5}}$, $\Delta_{\nu}=0.124$,
$\Delta_{\tau_{Y}}=3.84\times 10^{3}$ $\Delta_{\tau_{Y}}=3.83\times 10^{3}$ $\Delta_{\tau_{Y}}=\mathbf{0.85\times 10^{3}}$
Trophy $\Delta_{\theta_{fric}}=29.48^{\circ}$ $\Delta_{\theta_{fric}}=29.95^{\circ}$ $\Delta_{\theta_{fric}}=\mathbf{0.5^{\circ}}$
Table 6: Ablation study of different losses for system identification performance on the synthetic dataset. $\Delta_{*}$ denotes the absolute error (AE) $\downarrow$ for the material parameter *.
@@ C.3 Additional Evaluation Using Single-view Input
To evaluate our approach, we perform experiments using videos generated by PhysGen and Kling [19], with input images sourced from the PhysGenBench [25] dataset. As shown in Fig. 8, the visual results highlight the capability of our method to capture realistic deformations and motion trajectories guided by videos from both PhysGen and Kling. These results demonstrate the adaptability and robustness of our method, even when limited to single-view inputs, effectively simulating complex interactions and maintaining high fidelity across varying scenarios.
@@ C.4 Ablation Study
Effectiveness of Optical Flow Guidance: We provide detailed ablation study results of each object in Tab. 6 and qualitative comparisons in Fig. 6. Our method consistently achieves the lowest absolute error (AE) across a range of material parameters compared to other loss functions. The simulations generated by our approach closely align with the ground truth (GT), both in terms of material deformation and shape fidelity. These results show that optical flow guidance effectively captures complex material behaviors and enables precise material parameter optimization.

### v2

@@ 5 Conclusion
Acknowledgement: The work described in this paper was supported by a grant from the Research Grants Council of the Hong Kong Special Administrative Region, China (Project No. GRF/PolyU 15606321).
@@ B.1 Optimization Details
In our experiments, we implement the differentiable MPM using Warp [13]. Besides, we use RAFT [40] to compute optical flow, which serves as a critical component for guiding the optimization process. For the real-world dataset, we segment the foreground points where input forces are applied using SAM2Point [12], ensuring precise identification of the regions impacted by force. Specific forces and durations are detailed in Tab. 7. Video generation is facilitated by CogVideoX [46], which leverages a text prompt (refer to Tab. 6) combined with a selected frame from the dataset to guide the synthesis process. All optimization tasks are performed on a single NVIDIA A800 GPU.
@@ B.2 Baselines
PhysGaussian/PhysDreamer/Physics3D: PhysGaussian utilizes a fixed set of manually defined material properties, including material type, density, and associated parameters, without further optimization. This setup is comparable to the conditions outlined in Sec. 4.3 (1). Changes in material type, density, or further parameter optimization can lead to different simulation outcomes. To highlight the benefits of incorporating a large pre-trained visual foundation model for physics reasoning and parameter refinement, we apply the same input force to each scene and compare our results with PhysGaussian. While PhysDreamer and Physics3D include methods for parameter optimization, they are limited to the elastic material type. For consistency and fair comparison, we assign elastic as the material type for PhysDreamer and elastic with viscoelasticity for Physics3D. Besides, in our experiments, we apply larger forces (refer to Tab. 7) than previous baselines on the real-world dataset. The weaker performance of baselines under these conditions highlights our approach’s advantage in leveraging GPT initialization and optical flow guidance.
@@
Scene Force (s)  ($(x,y,z)$ for most cases) Duration (s)
Alocasia $(0.25,0,0)$ 1
Carnation $(-0.1,0,0)$ 1
Hat $(1,-2,1)$ 1
Telephone $(-1,0,0)$ 2
Fox $(0,-0.5,0.25)\xrightarrow{}(0,0,-0.5)\xrightarrow{}(0,0.5,0.25)$ $\xrightarrow{}$$\xrightarrow{}$1
Plane $rotation\ scale=-10\xrightarrow{}-5$ $\xrightarrow{}$1
Kitchen $(0,0,0.1)$ 1
Jam $(0.2,0,0)\xrightarrow{}(0.1,0.2,0)$ $\xrightarrow{}$1
Sandcastle $release\ n_{layer}=50$
Table 7: Simulation forces and durations on real-world dataset.
@@ C.1 Human Evaluation
C.2 Additional Experimental Results of Synthetic and Real-world Datasets
We provide additional comparisons between our method and baselines on both synthetic (see Tab. 9 and Fig. 9) and real-world (see Fig. 11) datasets. These results emphasize the effectiveness of our method in handling diverse material behaviors and complex scenarios. To further enhance understanding and visualization, we offer more results in videos accessible through our project page: https://zhuomanliu.github.io/PhysFlow
         PhysGen          PhysGaussian          Ours
         0.54          2.95          0.85
[…]
@@ C.3 Additional Evaluation Using Single-view Input
To evaluate our approach, we perform experiments using videos generated by PhysGen and Kling [20], with input images sourced from the PhyGenBench [29] dataset. As shown in Fig. 10, the visual results highlight the capability of our method to capture realistic deformations and motion trajectories guided by videos from both PhysGen and Kling. We also report the quantitative results (ECMS) on PhysGen scenes in Tab. 8. These results demonstrate the adaptability and robustness of our method, even when limited to single-view inputs, effectively simulating complex interactions and maintaining high fidelity across varying scenarios.
@@ C.4 Ablation Study
Effectiveness of Optical Flow Guidance: We performed ablations of $\mathcal{L}_{flow}$ on the entire synthetic dataset, which includes 9 cases across various material types. We provide detailed ablation results of each object in Tab. 10 and qualitative comparisons in Fig. 8. Our method consistently achieves the lowest relative error (RE) and ECMS across a range of material parameters compared to other loss functions. The simulations generated by our approach closely align with the ground-truth, both in terms of material deformation and shape fidelity. These results show that optical flow guidance effectively captures complex material behaviors and enables precise material parameter optimization.
Effectiveness of GPT Initialization: Along with experimental results in Sec. 3 and Fig. 7, GPT initialization reduces errors, and our full method with optical flow guidance achieves the lowest ECMS score and realistic motion.
[…]
@@ D Limitations
      Object       $\mathcal{L}_{render}$       $\mathcal{L}_{sds}$       Ours ($\mathcal{L}_{flow}$)
      Droplet       $\delta_{\mu}=0.045$, $\delta_{\kappa}=0.080$       $\delta_{\mu}=\mathbf{0.005}$, $\delta_{\kappa}=0.820$       $\delta_{\mu}=0.004$, $\delta_{\kappa}=\mathbf{0.030}$
      Letter       $\delta_{\mu}=0.162$, $\delta_{\kappa}=0.350$       $\delta_{\mu}=0.050$, $\delta_{\kappa}=\mathbf{0.000}$       $\delta_{\mu}=\mathbf{0.023}$, $\delta_{\kappa}=0.893$
      Cream       $\delta_{\mu}=11.100$, $\delta_{\kappa}=0.570$,       $\delta_{\mu}=\mathbf{0.030}$, $\delta_{\kappa}=\mathbf{0.480}$,       $\delta_{\mu}=\mathbf{0.080}$, $\delta_{\kappa}=0.755$,
      $\delta_{\tau_{Y}}=0.053$, $\delta_{\eta}=0.440$       $\delta_{\tau_{Y}}=\mathbf{0.007}$, $\delta_{\eta}=\mathbf{0.340}$       $\delta_{\tau_{Y}}=0.577$, $\delta_{\eta}=0.750$
      Toothpaste       $\delta_{\mu}=0.302$, $\delta_{\kappa}=1.220$,       $\delta_{\mu}=0.162$, $\delta_{\kappa}=\mathbf{0.076}$,       $\delta_{\mu}=\mathbf{0.015}$, $\delta_{\kappa}=0.136$,
      $\delta_{\tau_{Y}}=0.140$, $\delta_{\eta}=\mathbf{0.023}$       $\delta_{\tau_{Y}}=\mathbf{0.130}$, $\delta_{\eta}=0.090$       $\delta_{\tau_{Y}}=0.232$, $\delta_{\eta}=0.539$
      Torus       $\delta_{E}=0.040$, $\delta_{\nu}=0.073$       $\delta_{E}=\mathbf{0.010}$, $\delta_{\nu}=\mathbf{0.017}$       $\delta_{E}=0.039$, $\delta_{\nu}=0.989$
      Bird       $\delta_{E}=0.073$, $\delta_{\nu}=0.090$       $\delta_{E}=\mathbf{0.027}$, $\delta_{\nu}=\mathbf{0.053}$       $\delta_{E}=0.040$, $\delta_{\nu}=0.571$
      Playdoh       $\delta_{E}=0.920$, $\delta_{\nu}=0.093$,       $\delta_{E}=0.210$, $\delta_{\nu}=\mathbf{0.073}$,       $\delta_{E}=\mathbf{0.104}$, $\delta_{\nu}=0.327$,
      $\delta_{\tau_{Y}}=0.097$       $\delta_{\tau_{Y}}=\mathbf{0.013}$       $\delta_{\tau_{Y}}=0.027$
      Cat       $\delta_{E}=0.839$, $\delta_{\nu}=0.023$,       $\delta_{E}=\mathbf{0.020}$, $\delta_{\nu}=\mathbf{0.013}$,       $\delta_{E}=0.387$, $\delta_{\nu}=0.414$,
      $\delta_{\tau_{Y}}=0.073$       $\delta_{\tau_{Y}}=\mathbf{0.023}$       $\delta_{\tau_{Y}}=0.221$
      Trophy             $\delta_{\theta_{fric}}=0.098$       $\delta_{\theta_{fric}}=0.117$       $\delta_{\theta_{fric}}=\mathbf{0.013}$
Table 9: Comparisons with baselines for system identification performance on the synthetic dataset. $\delta_{*}$ denotes the relative error (RE) $\downarrow$ for the material parameter *.
      Object       $\mathcal{L}_{render}$       $\mathcal{L}_{sds}$       Ours ($\mathcal{L}_{flow}$)
      Droplet       $\delta_{\mu}=0.230$, $\delta_{\kappa}=0.731$       $\delta_{\mu}=1.515$, $\delta_{\kappa}=0.449$       $\delta_{\mu}=\mathbf{0.004}$, $\delta_{\kappa}=\mathbf{0.030}$
      Letter       $\delta_{\mu}=0.250$, $\delta_{\kappa}=0.918$       $\delta_{\mu}=0.575$, $\delta_{\kappa}=\mathbf{0.827}$       $\delta_{\mu}=\mathbf{0.023}$, $\delta_{\kappa}=0.893$
      Cream       $\delta_{\mu}=0.160$, $\delta_{\kappa}=0.732$,       $\delta_{\mu}=3.320$, $\delta_{\kappa}=\mathbf{0.004}$,       $\delta_{\mu}=\mathbf{0.080}$, $\delta_{\kappa}=0.755$,
      $\delta_{\tau_{Y}}=\mathbf{0.070}$, $\delta_{\eta}=0.841$       $\delta_{\tau_{Y}}=0.843$, $\delta_{\eta}=0.893$       $\delta_{\tau_{Y}}=0.577$, $\delta_{\eta}=\mathbf{0.750}$
      Toothpaste       $\delta_{\mu}=0.283$, $\delta_{\kappa}=0.173$,       $\delta_{\mu}=0.255$, $\delta_{\kappa}=0.141$,       $\delta_{\mu}=\mathbf{0.015}$, $\delta_{\kappa}=\mathbf{0.136}$,
      $\delta_{\tau_{Y}}=0.480$, $\delta_{\eta}=0.249$       $\delta_{\tau_{Y}}=0.375$, $\delta_{\eta}=\mathbf{0.178}$       $\delta_{\tau_{Y}}=\mathbf{0.232}$, $\delta_{\eta}=0.539$
      Torus             $\delta_{E}=0.237$, $\delta_{\nu}=\mathbf{0.586}$       $\delta_{E}=0.369$, $\delta_{\nu}=0.771$       $\delta_{E}=\mathbf{0.039}$, $\delta_{\nu}=0.989$
      Bird       $\delta_{E}=0.120$, $\delta_{\nu}=0.616$       $\delta_{E}=0.797$, $\delta_{\nu}=0.727$       $\delta_{E}=\mathbf{0.040}$, $\delta_{\nu}=\mathbf{0.571}$
      Playdoh       $\delta_{E}=0.483$, $\delta_{\nu}=0.196$,       $\delta_{E}=0.828$, $\delta_{\nu}=\mathbf{0.165}$,       $\delta_{E}=\mathbf{0.104}$, $\delta_{\nu}=0.327$,
      $\delta_{\tau_{Y}}=0.942$       $\delta_{\tau_{Y}}=0.943$       $\delta_{\tau_{Y}}=\mathbf{0.027}$
      Cat       $\delta_{E}=\mathbf{0.167}$, $\delta_{\nu}=\mathbf{0.290}$,       $\delta_{E}=0.644$, $\delta_{\nu}=0.623$,       $\delta_{E}=0.387$, $\delta_{\nu}=0.414$,
      $\delta_{\tau_{Y}}=0.652$       $\delta_{\tau_{Y}}=1.288$       $\delta_{\tau_{Y}}=\mathbf{0.221}$
      Trophy       $\delta_{\theta_{fric}}=0.173$       $\delta_{\theta_{fric}}=0.305$       $\delta_{\theta_{fric}}=\mathbf{0.013}$
Table 10: Ablation study of different losses for system identification performance on the synthetic dataset. $\delta_{*}$ denotes the relative error (RE) $\downarrow$ for the material parameter *.